# Linear combinations of cluster mean-field states applied to spin systems

### v1

@@ IV Correlation and non-singlet cases
The methodology presented in the previous section limits us to spin systems which have $S^{2}$ as a symmetry. In this section, we extend this methodology to systems which do not have this symmetry, and simultaneously incorporate correlations within the cMF scheme, similarly to what is done for cluster-CI (cCI).Abraham and Mayhall (2020)
IV.1 Dimer case
@@
In the dimer case, our approach does not involve optimizing the coefficients of an unrestricted cMF wavefunction. Instead, we have a linear combination of RcMF states. When the Hamiltonian has $S^{2}$ symmetry, we can include correlations by exciting one or more tile from the singlet to the triplet. For Hamiltonians which do not have $S^{2}$ symmetry, such as the XXZ model, this procedure is needed so that the wave function does not have a symmetry it should not possess. For a given tiling with $N$ 2-site tiles, this gives us $2^{N}$ states. In practice we will generally wish to limit the number of such excitations. In addition to the RcMF state, we will include all cases in which one or two tiles are in the triplet configuration, which we refer to as “single” and “double” excitations, respectively. This follows the usual terminology for cluster-based CI where a single excitation has one cluster not in its ground state, a double excitation has two clusters not in the ground state, and so on.
@@
We may choose to include higher excitations, but if we do, we restrict the excited clusters to all be adjacent to one another, similarly to Ref. 33. We symbolize this truncation scheme like dLC-cMF[n]-exc, where $d$ refers to the bond level and $n$ refers to the cluster size as noted earlier, and $exc$ refers to the triplet states. For example, 3LC-cMF[2]-sd refers to a computation utilizing all dimer coverings with bond level of 3, with all the one-at-a-time and two-at-a-time triplet cases. Lastly, we note that for Hamiltonians with $S^{2}$ symmetry, e.g. the $J_{1}-J_{2}$ Heisenberg model, we do not require single excitations as they inevitably change the $S^{2}$ quantum number.
@@
To demonstrate the efficiency of this correlation and truncation scheme, we present results for the 16-site XXZ Heiseinberg chain with OBC in Fig.6. As expected, including only single and double excitations yields results virtually indistinguishable from including all excitations. This result is not exact because each tile is still forced to have $S_{z}=0$, and the case where a given tile has $S_{z}=\pm 1$ has been totally excluded. These extra tilings would be included in a GcMFPapastathopoulos-Katsaros, Henderson, and Scuseria (2023) version of the theory, which we elect not to pursue in the present work. Nevertheless, this outcome is expected since the crucial part of the wavefunction arises from the various tiling schemes, with triplet excitations introducing secondary effects. Additionally, given the relatively small system size, we do not exclude many excitations. As a result, for much larger systems, we opt to include up to quadruple excitations, but triple and quadruple excitations are allowed only for contiguous dimers, as we have already noted. Our results for the 36-site $XXZ$ chain with OBC are summarized in Fig.7. Here, we can achieve an accurate approximation to the exact solution, especially for values of $\Delta>0$, despite the small bond length. The discrepancies around $\Delta=-1$ are not a significant concern, as our previous work Papastathopoulos-Katsaros, Henderson, and Scuseria (2023) has demonstrated how to obtain highly accurate energy estimates in that regime. A straightforward linear combination of GcMF with LC-cMF is likely to provide an excellent solution, which we can explore in future research. Lastly, we note that it is also possible to imagine a linear combination of dimer tilings with each dimer treated as UcMF, assuming that the Hamiltonian does not have $S^{2}$ symmetry.
@@
Figure 7: Energy per site obtained in 1LC-cMF[2]-sdtq (a linear combination of RcMF with 2-site tiles in a single tiling scheme and up to four-at-a-time triplets), 2LC-cMF[2]-sdtq (a linear combination of RcMF with 2-site tiles in 18 different tiling schemes and up to four-at-a-time triplets) and FCI for the 36-site $XXZ$ Heisenberg chain.
IV.2 Tetramer case
In the present section, we delve into the exploration of linear combinations involving larger clusters, specifically focusing on tetramers due to their viability for the systems we are examining. To the best of our knowledge, there is currently no readily available basis for tetramer coverings in closed form. However, this is not a significant constraint since our methods are designed with a truncation approach in mind. Previous researchJiménez-Hoyos and Scuseria (2015) has demonstrated that compact shapes generally yield more accurate estimates of the ground state energy. For this reason, we limit our consideration to 5 tetramer tiles, akin to those depicted in Fig. 8, which constitute the fundamental TetrisPajitnov (1985) pieces. Each of these 5 tetramer configurations employs a single tile type, albeit with variations in inversion, rotation, and reflection. While it is possible to conceive shapes that, when considering the spin configurations of the exact ground state for each Hamiltonian, might yield lower energies, our objective is to establish a method that operates in a black-box manner, and tiles with intricate shapes can pose challenges.
Figure 8: The 5 tetramer coverings used for 2D square lattices.
@@ IV.2 Tetramer case
We decide to simultaneously optimize the coefficients $c^{ijkl}$ for every tetramer covering, eliminating the need for additional cases like the dimers. In practical terms, if, for instance, we work with five tetramer coverings, we need to conduct a relatively modest $5\times 5$ non-orthogonal configuration interaction (NOCI) during each optimization step. Consequently, this optimization process is generally more challenging in terms of convergence and computational resources, although conceptually simpler.
@@
V Discussion
In section II, we explored the cluster mean-field method for addressing strongly-correlated spin systems. Building upon this foundation, in section III we introduced the linear combinations of cluster mean-field formalism for singlet ground states. In section IV, we extended this approach to arbitrary $S^{2}$, which further allowed us to incorporate additional correlations. Depending on the Hamiltonian, these extensions can be applied as variational approaches for the ground state wavefunction, providing relatively accurate estimates for the energy. Our findings clearly illustrate that LC-cMF methods can qualitatively capture the essential aspects of the ground state physics in benchmark models and can serve as valuable references for more advanced correlated methods, with mild computational cost in an efficient code. In addition, they offer compelling evidence that our method can be employed for a semi-quantitative description of the $0.4\lessapprox J_{2}/J_{1}\lessapprox 0.6$ regime of the 2D $J_{1}-J_{2}$ model, which is a particularly demanding regime for established methodologies.
In our earlier work, we discussed the correlation of GcMF, as detailed in Ref. 3, with the use of Jastrow-like operators. These operators have been applied previously in the context of the antisymmetrized geminal power (AGP) wavefunction, as highlighted in references such as 67 and 68, showcasing promising outcomes. The inclusion of such operators has the potential to enhance the accuracy of our cluster-based wavefunctions and expand their applicability to more intricate systems. Lastly, we can use a similar approach to the few determinant approximation (FED), Rodríguez-Guzmán, Jiménez-Hoyos, and Scuseria (2014a); Bytautas et al. (2014); Rodríguez-Guzmán et al. (2013); Jiménez-Hoyos, Rodríguez-Guzmán, and Scuseria (2013); Rodríguez-Guzmán, Jiménez-Hoyos, and Scuseria (2014b); Rodríguez-Guzmán et al. (2012) but instead of determinants we can utilize excited cluster product states and optimize them in a NOCI manner, similarly to the present work.
@@ V Discussion
Acknowledgements.

### v2

@@ 4 Correlation and non-singlet cases
The methodology presented in the previous section limits us to spin systems which have $S^{2}$ as a symmetry, because each tile, and therefore the system as a whole, is an $S^{2}$ eigenstate with $S^{2}=0$. In this section, we extend this methodology to systems which do not have this symmetry, and simultaneously incorporate correlations within the cMF scheme, similarly to what is done for cluster-CI (cCI).7, 8, 9
4.1 Dimer case
@@
For non-singlet cases, we could consider a simple linear combination of UcMF states (as defined in Eqn. 7). This would require us to optimize the coefficients $c^{ij}_{\uparrow\downarrow}$ and $c^{ij}_{\downarrow\uparrow}$, in each tile, for each tiling scheme in the linear combination. To simplify the calculation and reduce the cost, we try a different approach. For a given tiling scheme, we include the RcMF state and a handful of excited states in which one or more clusters have been excited from the singlet configuration (Eqn. 5) to the triplet (Eqn. 6). For Hamiltonians which do not have $S^{2}$ symmetry, such as the XXZ model, this procedure is needed so that the wave function does not have a symmetry it should not possess. For a given tiling with $N$ 2-site tiles, this gives us $2^{N}$ states. In practice we will generally wish to limit the number of such excitations. In addition to the RcMF state, we will include all cases in which one or two tiles are in the triplet configuration, which we refer to as “single” and “double” excitations, respectively. This follows the usual terminology for cluster-based configuration interaction where a single excitation has one cluster not in its ground state, a double excitation has two clusters not in the ground state, and so on.
@@
We may choose to include higher excitations, but if we do, we restrict the excited clusters to all be adjacent to one another, similarly to Ref. 38. We symbolize this truncation scheme as dLC-cMF[n]-exc, where $d$ refers to the bond level and $n$ refers to the cluster size as noted earlier, and $exc$ refers to the triplet states. For example, 3LC-cMF[2]-sd refers to a computation utilizing all dimer coverings with bond level of 3, with all the one-at-a-time and two-at-a-time triplet cases. Lastly, we note that for Hamiltonians with $S^{2}$ symmetry, e.g. the $J_{1}-J_{2}$ Heisenberg model, we do not require single excitations as they inevitably change the $S^{2}$ quantum number.
@@
To demonstrate the efficiency of this correlation and truncation scheme, we present results for 1D XXZ Heisenberg systems with OBC in Figs.6 and 7. In Fig.6, our results for the 16-site XXZ Heiseinberg chain with OBC are shown. As expected, including only single and double excitations yields results virtually indistinguishable from including all excitations. This result is not exact because each tile is still forced to have $S_{z}=0$, and the case where a given tile has $S_{z}=\pm 1$ has been totally excluded. These extra tilings would be included in a GcMF3 version of the theory, which we elect not to pursue in the present work. Nevertheless, this outcome is expected since the crucial part of the wavefunction arises from the various tiling schemes, with triplet excitations introducing secondary effects. Additionally, given the relatively small system size, we do not exclude many excitations. As a result, for much larger systems, we opt to include up to quadruple excitations, but triple and quadruple excitations are allowed only for contiguous dimers, as we have already noted. Our results for the 36-site $XXZ$ chain with OBC are summarized in Fig.7. Here, we can achieve an accurate approximation to the exact solution, especially for values of $\Delta>0$, despite the small bond length. The discrepancies around $\Delta=-1$ are not a significant concern, as our previous work 3 has demonstrated how to obtain highly accurate energy estimates in that regime. It is possible to imagine a linear combination of UcMF or even GcMF states, which would come with greater variatonal flexibility and should yield superior answers albeit with a significantly greater computational cost. Alternatively, since a single GcMF suffices for the region for $\Delta<-1$, we could also just include this one GcMF state in our linear combination.
@@
Figure 7: Energy per site obtained in UcMF[2], 2LC-cMF[2]-sd (a linear combination of RcMF with 2-site tiles in 18 different tiling schemes and up to two-at-a-time triplets), 2LC-cMF[2]-sdtq (a linear combination of RcMF with 2-site tiles in 18 different tiling schemes and up to four-at-a-time triplets) and FCI for the 36-site $XXZ$ Heisenberg chain.
4.2 Tetramer case
In the present section, we delve into the exploration of linear combinations involving larger clusters, specifically focusing on tetramers due to their viability for the systems we are examining. To the best of our knowledge, there is currently no readily available basis for tetramer coverings in closed form. However, this is not a significant constraint since our methods are designed with a truncation approach in mind. Previous research1 has demonstrated that compact shapes generally yield more accurate estimates of the ground state energy. For this reason, we limit our consideration to 5 tetramer tiles, akin to those depicted in Fig. 8, which constitute the fundamental Tetris70 pieces. Each of these 5 tetramer configurations employs a single tile type, albeit with variations in inversion, rotation, and reflection. While it is possible to conceive shapes that, when considering the spin configurations of the exact ground state for each Hamiltonian, might yield lower energies, our objective is to establish a method that operates in a black-box manner, and tiles with intricate shapes can pose challenges.
@@ 4.2 Tetramer case
where dimers have only one singlet state, in tetramers there are two (as well as three triplets and a quintet). This means that even for a restricted LC-cMF we would have to optimize the linear combination of the two singlet states used in each tile. We therefore decided to optimize all six of the coefficients $c^{ijkl}$ for every tile in every tetramer covering simultaneously. This increases the variational flexibility in our LC-cMF significantly, although with concomitant increase in computational cost. We note, though, that the excitation-based scheme we introduced in the previous section could be employed for tetramers as well, except that where a dimer has only one excited state, a tetramer has five.
@@
5 Discussion
In section 2, we explored the cluster mean-field method for addressing strongly-correlated spin systems. Building upon this foundation, in section 3 we introduced the linear combinations of cluster mean-field formalism for singlet ground states. In section 4, we extended this approach to arbitrary $S^{2}$, which further allowed us to incorporate additional correlations. Depending on the Hamiltonian, these extensions can be applied as variational approaches for the ground state wavefunction, providing relatively accurate estimates for the energy. Our findings clearly illustrate that LC-cMF methods can qualitatively capture the essential aspects of the ground state physics in benchmark models and can serve as valuable references for more advanced correlated methods, with mild computational cost in an efficient code. In addition, they offer compelling evidence that our method can be employed for a semi-quantitative description of the $0.4\lessapprox J_{2}/J_{1}\lessapprox 0.6$ regime of the 2D $J_{1}-J_{2}$ model, which is a particularly demanding regime for established methodologies.
In our earlier work, we discussed the correlation of GcMF, as detailed in Ref. 3, with the use of Jastrow-like operators. These operators have been applied previously in the context of the antisymmetrized geminal power (AGP) wavefunction, as highlighted in references such as 71 and 72, showcasing promising outcomes. The inclusion of such operators has the potential to enhance the accuracy of our cluster-based wavefunctions and expand their applicability to more intricate systems. Lastly, we can use a similar approach to the few determinant approximation (FED), 73, 74, 75, 76, 77, 78 but instead of determinants we can utilize excited cluster product states and optimize them in a NOCI manner, similarly to the present work.
@@ 5 Discussion
Author Information
Corresponding Author
Gustavo E. Scuseria - Rice University, Houston, Texas 77005, USA; Email: guscus@rice.edu
{acknowledgement}